# Stabilizer Entanglement Distillation and Efficient Fault-Tolerant Encoder

### v1

@@ Abstract
Entanglement is essential for quantum information processing but is limited by noise. We address this by developing high-yield entanglement distillation protocols with several advancements. (1) We extend the $2$-to-$1$ recurrence entanglement distillation protocol to higher-rate $n$-to-$\left(n-1\right)$ protocols that can correct any single-qubit errors. These protocols are evaluated through numerical simulations focusing on fidelity and yield. We also outline a method to adapt any classical error-correcting code for entanglement distillation, where the code can correct both bit-flip and phase-flip errors by incorporating Hadamard gates. (2) We propose a constant-depth decoder for stabilizer codes that transforms logical states into physical ones using single-qubit measurements. This decoder is applied to entanglement distillation protocols, reducing circuit depth and enabling protocols derived from advanced quantum error-correcting codes. We demonstrate this by evaluating the circuit complexity for entanglement distillation protocols based on surface codes and quantum convolutional codes. (3) Our stabilizer entanglement distillation techniques advance quantum computing. We propose a fault-tolerant protocol for constant-depth encoding and decoding of arbitrary quantum states, applicable to quantum low-density parity-check (qLDPC) codes and surface codes. This protocol is feasible with state-of-the-art reconfigurable atom arrays and surpasses the limits of conventional logarithmic depth encoders. Overall, our study integrates stabilizer formalism, measurement-based quantum computing, and entanglement distillation, advancing both quantum communication and computing.
@@ I Introduction
Entanglement is essential for quantum communication [1, 2, 3, 4], quantum metrology [5, 6, 7, 8], and distributed quantum computing [9, 10, 11, 12, 13], but it is severely limited by noise [14, 15, 16, 17, 18]. Entanglement distillation [19, 20, 21, 22], or entanglement purification, is a process that transforms previously shared, less entangled Bell states into fewer, but higher-quality states through local operations and classical communication (LOCC). This technique has been explored across various physical platforms [23, 24, 25, 26, 27], demonstrated experimentally [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and extended to multi-qubit systems [39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. However, the basic $2$-to-$1$ protocols [19, 21] are inefficient, invariably sacrificing at least half of the input states, regardless of their initial fidelity. The hashing protocol [20, 49], while offering adaptive yields based on the input fidelity, requires an infinitely large input, limiting its practical use. This highlights the need for more efficient entanglement distillation protocols capable of generating higher yields from finite input states.
@@
Research on efficient distillation protocols has explored strategies such as high-dimensional systems [50, 51, 52, 53, 54, 55], entanglement-assisted methods [56, 57, 58, 59, 60], symmetry-based searches [61, 62], and heuristic optimizations [63, 64, 65, 66, 67, 68, 69]. A key approach utilizes quantum error correction [70, 71, 72, 73, 74, 75, 76], which provides reliable performance guarantees. Specifically, in the so-called “stabilizer protocol” of entanglement distillation [77], Alice and Bob start with $n$ shared noisy Bell states. As shown in Figure 1(a), each party independently measures $n-k$ stabilizer generators on their $n$ qubits, projecting them onto logical subspaces. Alice sends her stabilizer parities to Bob, who identifies an error syndrome and corrects the errors by applying appropriate Pauli operators. After error correction, both parties perform unitary decoding, the inverse of the unitary encoding circuit, to transform the $k$ logical Bell states into $k$ physical Bell states. Crucially, it was shown that the same can be accomplished by a direct unitary decoding applied to both Alice’s and Bob’s qubits of the $n$ noisy Bell states [70, 71], as illustrated in Figure 1(b). This is followed by each party performing $n-k$ single-qubit measurements to reveal stabilizer parities, Alice communicating her parities to Bob, and Bob performing a recovery operation to his $k$ unmeasured qubits, resulting in $k$ distilled physical Bell states. This approach makes initial stabilizer measurements unnecessary. However, stabilizer measurements exhibit a more compact structure than unitary decoding circuits, especially evident in quantum low-density parity-check (qLDPC) codes [78] and surface codes [79, 80]. These codes feature a constant number of qubits per stabilizer generator and a constant number of generators acting on each qubit, enabling constant-depth stabilizer measurements. The primary challenge lies in developing an efficient entanglement distillation scheme that leverages the compactness of stabilizer measurements.
@@
    We review stabilizer entanglement distillation and apply it to Bennett’s and Deutsch’s protocols [19, 21], mapping them onto a $\left[2,1,2\right]$ classical error detection code. We demonstrate that bilateral Hadamard gates on the distillation output can alternate Pauli-$X$ and Pauli-$Z$ errors while preserving the standard Bell states. Consequently, two sequential applications of these protocols can effectively address any single-qubit errors. Drawing on this insight, we extend the $2$-to-$1$ protocol to higher-rate $n$-to-$\left(n-1\right)$ protocols, corresponding to an error detection process using a $\left[\!\left[n^{2},\left(n-1\right)^{2},2\right]\!\right]$ quantum code. We evaluate the performance of these protocols through numerical simulations, assessing both input-output fidelity and yield. Additionally, we present a method to adapt any classical error correcting code to an entanglement distillation protocol using a two-iteration process. The first iteration corrects bit-flip errors, then Hadamard gates convert phase-flip errors to bit-flip errors, which the second iteration corrects. Since the errors in both iterations are classical, the error-correcting capability matches that of the classical code used.
@@
    General Code $\mathcal{O}\left(n\right)$ $\mathcal{O}\left(n\right)$
    Surface Code $\geq\mathcal{O}\left(\log n\right)$ $\mathcal{O}\left(1\right)$
@@
    We propose a fault-tolerant protocol for constant-depth encoding and decoding of arbitrary quantum states, applicable to qLDPC codes and surface codes. The protocol incorporates stabilizer measurements and efficient decoding to establish physical-logical Bell states. We enhance fault tolerance through entanglement distillation of these Bell states and employ quantum teleportation for encoding and decoding quantum states. The protocol is feasible with current reconfigurable atom arrays [88, 89] and surpasses the conventional depth limits for surface codes, lower bounded by $\mathcal{O}\left(\log\left(n\right)\right)$ [82, 83], where $n$ is the number of qubits.
@@ II.1 Error Model
signifies that an arbitrary operator $M$ applied to Alice’s qubit of the Bell state is equivalent to its transpose applied to Bob’s qubit. This allows us to simplify the error model by assuming that only Bob’s qubit is subjected to a noisy channel, while Alice’s qubit remains error-free. For our model, we consider a depolarizing channel that applies Pauli errors randomly with equal probability. The resulting noisy Bell state is given by
@@
where $\left|\Phi_{n}^{+}\right\rangle=\left|\Phi^{+}\right\rangle^{\otimes n}$ denotes the tensor product of $n$ standard Bell state. Here, $I$ is the identity operator acting on Alice’s qubits, and $E$ represents an $n$-qubit Pauli operator, expressed as $E=E_{1}\otimes E_{2}\otimes\cdots\otimes E_{n}$, with each $E_{i}\in\{I,X,Y,Z\}$ acting on Bob’s qubits indexed from $1$ to $n$. The occurrence of $E=I^{\otimes n}$ indicates an error-free event, while $E_{i}\neq I$ for any $i$ signifies an error on the $i$-th qubit, marking an erroneous event. For brevity, we omit the $\otimes$ symbol for one-partite tensor products in subsequent discussions, for example, writing $E=E_{1}E_{2}\cdots E_{n}$.
@@ II.2 Entanglement Distillation Protocol
To facilitate entanglement distillation, we use a stabilizer code parameterized by $\left[\!\left[n,k,d\right]\!\right]$ [90]. Here, $n$ represents the number of physical qubits, $k$ the number of logical qubits encoded, and $d$ the minimum distance between logical qubits. The code features $n-k$ stabilizer generators, which are essentially $n$-qubit Pauli operators, each denoted as $g_{i}$ for $i=1,2,\ldots,n-k$. These generators are independent and mutually commuting. If a state $\left|\psi\right\rangle$ is an eigenvector of the generator $g_{i}$, satisfying
@@
    Alice performs $n-k$ stabilizer measurements on her qubits. These measurements project the state onto $P_{i}^{\left(s_{i}\right)}\otimes E\left|\Phi_{n}^{+}\right\rangle$, where $P_{i}^{\left(s_{i}\right)}$ is the projector corresponding to the measurement outcome $s_{i}$. Utilizing operator commutation, projector property $P^{2}=P$, and Bell-state matrix identity, the post-measurement state can be represented as
@@
    Thus, Alice’s measurements project both her qubits and Bob’s qubits onto subspaces stabilized by generators $g_{i}$ and $g_{i}^{T}$, respectively. If $g_{i}$ contains an odd number of Pauli-$Y$s, then $g_{i}^{T}=-g_{i}$; otherwise, $g_{i}^{T}=g_{i}$.
@@
    Alice communicates her measurements, $\bm{s}_{A}=s_{1}s_{2}\ldots s_{n-k}$, to Bob. This allows Bob to know the parities of his qubits before any errors. The parities are given by $\bm{s}_{B}=s_{1}^{\prime}s_{2}^{\prime}\ldots s_{n-k}^{\prime}$. If a generator $g_{i}$ contains an odd number of Pauli-$Y$s, then $s_{i}^{\prime}=s_{i}\oplus 1$; otherwise, $s_{i}^{\prime}=s_{i}$.
@@ III Application I: Extended Recurrence Entanglement Distillation Protocol
We apply the stabilizer formalism to the well-known 2-to-1 recurrence entanglement distillation. Bennett’s protocol [19] can be seen as a stabilizer entanglement distillation in error detection mode. The code uses a single generator, $X_{1}X_{2}$, which detects single-qubit $Y$ and $Z$ errors but not $X$ errors. To address $X$ errors, Bennett et al. used twirling to randomize them, enabling detection in later iterations. Alternatively, Alice and Bob can apply Hadamard gates to their output qubits. The bilateral Hadamard gates preserve the standard Bell state and switch $X$ and $Z$ errors, as
@@
This allows the remaining $Z$ errors to be addressed in the next iteration, which is more efficient than twirling. The revised protocol is equivalent to Deutsch’s protocol [21], where the stabilizer code $Y_{1}Y_{2}$ is used to detect $X$ and $Z$ errors, and $R_{x}\left(\pi/2\right)$ is used to switch $Y$ and $Z$ errors. Since this protocol can detect $X$, $Y$, and $Z$ errors over two sequential iterations, it can detect any single-qubit error according to the discretization of quantum errors [90]. This explains why the recurrence protocols are effective in various error models, not just the Werner-state model.
Based on the above insight, we can extend the $2$-to-$1$ recurrence protocol to a higher-rate $n$-to-$\left(n-1\right)$ protocol using a stabilizer code $X_{1}X_{2}\cdots X_{n}$. For example, consider the $3$-to-$2$ entanglement distillation circuit illustrated in Figure  2, constructed using the unitary decoding method. The two dashed boxes represent circuits for two sequential iterations. In each box, the projective measurement on the first qubit reveals the parity of $X_{1}X_{2}X_{3}$. By comparing parities, Alice and Bob can detect single-qubit $Y$ or $Z$ errors. Hadamard gates on the remaining qubits convert $X$ errors to $Z$ errors, which are detected in the next iteration. The protocol is similar to a Calderbank-Shor-Steane (CSS) construction [91, 92], with one code addressing bit-flip errors and the other addressing phase-flip errors. By permuting the measurements and gates in Figure 2, we obtain the stabilizers for two iterations:
@@
which characterize a $\left[\!\left[9,4,2\right]\!\right]$ quantum code capable of detecting a single-qubit error. For the $n$-to-$\left(n-1\right)$ protocol, two iterations correspond to a $\left[\!\left[n^{2},\left(n-1\right)^{2},2\right]\!\right]$ quantum code.
To prevent correlated multi-qubit errors, it is essential to use outputs from different iterations as inputs in subsequent iterations. For example, in Figure 2, an $X_{1}$ error on the control qubit can propagate through the CNOT gates to become $X_{1}X_{2}X_{3}$. Assuming a depolarizing channel with an error probability of $p$ per input qubit, the effective error probability in the first iteration is $\frac{2p}{3}$ due to $Y$ and $Z$ errors. In the second iteration, the effective error probability is also $\frac{2p}{3}$ due to undetected $X$ errors on the output and control qubits. Thus, if the input meets the fidelity threshold initially, it will meet the threshold in subsequent iterations. This result applies to other $n$-to-$\left(n-1\right)$ recurrence protocols as well.
@@
where $k$ and $n$ are the logical and physical qubits in an $\left[\!\left[n,k,d\right]\!\right]$ code, and $P_{S}$ is the success probability. Our simulation uses the QuantumClifford.jl package, which utilizes the stabilizer tableaux formalism [93]. Each protocol undergoes two iterations, corresponding to an $n^{2}$-to-$\left(n-1\right)^{2}$ entanglement distillation, capable of addressing any single-qubit errors. We use Werner states as inputs, which are mixtures of a standard Bell state and a completely mixed state, as described by Equation 2. We assume error-free local quantum operations and classical communication (LOCC), a valid assumption given that channel noise significantly exceeds local noise.
@@
Figure 3(a) shows the fidelity of distilled Bell states as a function of input fidelity for several distillation protocols. Circles mark the threshold where output fidelity exceeds input fidelity, which increases with the code rate. Figure 3(b) shows yield as a function of input fidelity. The $4$-to-$1$ protocol has the highest yield for input fidelities from $0.5$ to $0.887$. Beyond this range, higher-rate protocols are superior. Near ideal fidelity, yield is determined by code rate. Achieving optimal yield at high fidelity requires large $n$ in $n^{2}$-to-$\left(n-1\right)^{2}$ protocols.
@@
We can further extend the recurrence protocol by adapting classical codes. In the first iteration, the code corrects bit-flip errors. Hadamard gates then convert the remaining phase-flip errors to bit-flip errors, which the code can correct in the second iteration. Because both iterations involve classical errors, the error-correcting capabilities match those of the classical codes used. Moreover, different codes can be employed in each iteration, allowing for customization based on the strength of phase-flip and bit-flip errors. In practice, noisy Bell states can be non-Werner [94, 95, 96]. Combining the stabilizer formalism with fidelity and yield aids in designing effective distillation protocols.
IV Decoding Stabilizer Code by Single-Qubit Measurements

### v2

@@ Abstract
Entanglement is essential for quantum information processing, but is limited by noise. We address this by developing high-yield entanglement distillation protocols with several advancements. (1) We extend the $2$-to-$1$ recurrence entanglement distillation protocol to higher-rate $n$-to-$\left(n-1\right)$ protocols that can correct any single-qubit errors. These protocols are evaluated through numerical simulations focusing on fidelity and yield. We also outline a method to adapt any classical error-correcting code for entanglement distillation, where the code can correct both bit-flip and phase-flip errors by incorporating Hadamard gates. (2) We propose a constant-depth decoder for stabilizer codes that transforms logical states into physical ones using single-qubit measurements. This decoder is applied to entanglement distillation protocols, reducing circuit depth and enabling protocols derived from high-performance quantum error-correcting codes. We demonstrate this by evaluating the circuit complexity for entanglement distillation protocols based on surface codes and quantum convolutional codes. (3) Our stabilizer entanglement distillation techniques advance quantum computing. We propose a fault-tolerant protocol for constant-depth encoding and decoding of arbitrary states in surface codes, with potential extensions to more general quantum low-density parity-check codes. This protocol is feasible with state-of-the-art reconfigurable atom arrays and surpasses the limits of conventional logarithmic depth encoders. Overall, our study integrates stabilizer formalism, measurement-based quantum computing, and entanglement distillation, advancing both quantum communication and computing.
@@ I Introduction
Entanglement is essential for quantum communication [1, 2, 3, 4], quantum metrology [5, 6, 7, 8], and distributed quantum computing [9, 10, 11, 12, 13], but it is severely limited by noise [14, 15, 16, 17, 18]. Entanglement distillation [19, 20, 21, 22], or entanglement purification, is a process that transforms previously shared, less entangled Bell states into fewer, but higher-quality states through local operations and classical communication. This technique has been explored across various physical platforms [23, 24, 25, 26, 27], demonstrated experimentally [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and extended to multiqubit systems [39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. However, the basic $2$-to-$1$ protocols [19, 21] are inefficient, invariably sacrificing at least half of the input states, regardless of their initial fidelity. The hashing protocol [20, 49], while offering adaptive yields based on the input fidelity, requires an infinitely large input, limiting its practical use. This highlights the need for more efficient entanglement distillation protocols capable of generating higher yields from finite input states.
@@
Research on efficient distillation protocols has explored strategies such as high-dimensional systems [50, 51, 52, 53, 54, 55], entanglement-assisted methods [56, 57, 58, 59, 60], symmetry-based searches [61, 62], and heuristic optimizations [63, 64, 65, 66, 67, 68, 69]. A key approach utilizes quantum error correction [70, 71, 72, 73, 74, 75, 76], which provides reliable performance guarantees. Specifically, in the so-called “stabilizer protocol” of entanglement distillation [77], Alice and Bob start with $n$ shared noisy Bell states. As shown in Figure 1(a), each party independently measures $n-k$ stabilizer generators on their $n$ qubits, projecting them onto logical subspaces. Alice sends her stabilizer parities to Bob, who identifies an error syndrome and corrects the errors by applying appropriate Pauli operators. After error correction, both parties perform unitary decoding, the inverse of the unitary encoding circuit, to transform the $k$ logical Bell states into $k$ physical Bell states. Crucially, it was shown that the same can be accomplished by a direct unitary decoding applied to both Alice’s and Bob’s qubits of the $n$ noisy Bell states [70, 71], as illustrated in Figure 1(b). This is followed by each party performing $n-k$ single-qubit measurements to reveal stabilizer parities, Alice communicating her parities to Bob, and Bob performing a recovery operation to his $k$ unmeasured qubits, resulting in $k$ distilled physical Bell states. This approach makes initial stabilizer measurements unnecessary. However, stabilizer measurements exhibit a more compact structure than unitary decoding circuits, particularly in quantum low-density parity-check (QLDPC) codes [78], with surface codes being a prominent example  [79, 80, 81]. These codes feature a constant number of qubits per stabilizer generator and a constant number of generators acting on each qubit, enabling constant-depth stabilizer measurements. The primary challenge lies in developing an efficient entanglement distillation scheme that leverages the compactness of stabilizer measurements.
@@
    We review stabilizer entanglement distillation and apply it to the protocols of Bennett et al. [19] and Deutsch et al. [21], mapping these protocols onto a $\left[2,1,2\right]$ classical error detection code. We demonstrate that bilateral Hadamard gates on the distillation output can alternate Pauli-$X$ and Pauli-$Z$ errors while preserving the standard Bell states. Consequently, two sequential applications of these protocols can effectively address any single-qubit errors. Drawing on this insight, we extend the $2$-to-$1$ protocol to higher-rate $n$-to-$\left(n-1\right)$ protocols, corresponding to an error detection process using a $\left[\!\left[n^{2},\left(n-1\right)^{2},2\right]\!\right]$ quantum code. We evaluate the performance of these protocols through numerical simulations, assessing both input-output fidelity and yield. Additionally, we present a method to adapt any classical error-correcting code to an entanglement distillation protocol using a two-iteration process. The first iteration corrects bit-flip errors, then Hadamard gates convert phase-flip errors to bit-flip errors, which the second iteration corrects. Since the errors in both iterations are classical, the error-correcting capability matches that of the classical code used.
@@
    General Code $O\left(n\right)$ $O\left(n\right)$
    Surface Code $\Omega\left(\log n\right)$ $O\left(1\right)$
@@
    We propose a fault-tolerant protocol for constant-depth encoding and decoding of arbitrary quantum states, applicable to QLDPC codes and surface codes. The protocol incorporates stabilizer measurements and efficient decoding to establish physical-logical Bell states. We enhance fault tolerance through entanglement distillation of these Bell states and employ quantum teleportation for encoding and decoding quantum states. The protocol is feasible with current reconfigurable atom arrays [89, 90] and surpasses the conventional depth limits for surface codes, lower bounded by $\Omega\left(\log\left(n\right)\right)$ [83, 84], where $n$ is the number of qubits.
@@ II.1 Error Model
signifies that an arbitrary operator $M$ applied to Alice’s qubit of the standard Bell state $\left|\Phi^{+}\right\rangle$ is equivalent to its transpose applied to Bob’s qubit. This allows us to simplify the error model by assuming that only Bob’s qubit is subjected to a noisy channel, while Alice’s qubit remains error-free (see Appendix A for a detailed calculation). For our model, we consider a depolarizing channel that applies Pauli errors randomly with equal probability. The resulting noisy Bell state is given by
@@
where $\left|\Phi_{n}^{+}\right\rangle=\left|\Phi^{+}\right\rangle^{\otimes n}$ denotes the tensor product of $n$ standard Bell states. Here, $I$ is the identity operator acting on Alice’s qubits, and $E$ represents an $n$-qubit Pauli operator, expressed as $E=E_{1}\otimes E_{2}\otimes\cdots\otimes E_{n}$, with each $E_{i}\in\{I,X,Y,Z\}$ acting on Bob’s qubits indexed from $1$ to $n$. The occurrence of $E=I^{\otimes n}$ indicates an error-free event, while $E_{i}\neq I$ for any $i$ signifies an error on the $i$-th qubit, marking an erroneous event. For brevity, we omit the $\otimes$ symbol for one-partite tensor products in subsequent discussions, for example, writing $E=E_{1}E_{2}\cdots E_{n}$.
@@ II.2 Entanglement Distillation Protocol
To facilitate entanglement distillation, we use a stabilizer code parameterized by $\left[\!\left[n,k,d\right]\!\right]$ [91]. Here, $n$ represents the number of physical qubits, $k$ the number of logical qubits encoded, and $d$ the code distance, defined as the minimum weight of any logical operator. The code features $n-k$ stabilizer generators, which are essentially $n$-qubit Pauli operators, each denoted as $g_{i}$ for $i=1,2,\ldots,n-k$. These generators are independent and mutually commuting. If a state $\left|\psi\right\rangle$ is an eigenvector of the generator $g_{i}$, satisfying
@@
    Alice performs $n-k$ stabilizer measurements on her qubits. These measurements project the state onto $P_{i}^{\left(s_{i}\right)}\otimes E\left|\Phi_{n}^{+}\right\rangle$, where $P_{i}^{\left(s_{i}\right)}$ is the projector corresponding to the measurement outcome $s_{i}$. Utilizing operator commutation, projector property $P^{2}=P$, and Bell-state matrix identity associated with the standard Bell state $\left|\Phi^{+}\right\rangle$, the postmeasurement state can be represented as
@@
    Thus, Alice’s measurements project both her qubits and Bob’s qubits onto subspaces stabilized by generators $g_{i}$ and $g_{i}^{T}$, respectively. If $g_{i}$ contains an odd number of Pauli-$Y$ operators, then $g_{i}^{T}=-g_{i}$; otherwise, $g_{i}^{T}=g_{i}$.
@@
    Alice communicates her measurements, $\bm{s}_{A}=s_{1}s_{2}\ldots s_{n-k}$, to Bob. This allows Bob to know the parities of his qubits before any errors. The parities are given by $\bm{s}_{B}=s_{1}^{\prime}s_{2}^{\prime}\ldots s_{n-k}^{\prime}$. If a generator $g_{i}$ contains an odd number of Pauli-$Y$ operators, then $s_{i}^{\prime}=s_{i}\oplus 1$; otherwise, $s_{i}^{\prime}=s_{i}$.
@@ III Application I: Extended Recurrence Entanglement Distillation Protocol
We apply the stabilizer formalism to the well-known 2-to-1 recurrence entanglement distillation. The protocol by Bennett et al. [19] can be seen as a stabilizer entanglement distillation in error detection mode. The code uses a single generator, $X_{1}X_{2}$, which detects single-qubit $Y$ and $Z$ errors but not $X$ errors. To address $X$ errors, Bennett et al. used twirling to randomize them, enabling detection in later iterations. Alternatively, Alice and Bob can apply Hadamard gates to their output qubits. The bilateral Hadamard gates preserve the standard Bell state and switch $X$ and $Z$ errors, as
@@
This allows the remaining $Z$ errors to be addressed in the next iteration, which is more efficient than twirling. The revised protocol is equivalent to the protocol by Deutsch et al. [21], where the stabilizer code $Y_{1}Y_{2}$ is used to detect $X$ and $Z$ errors, and $R_{x}\left(\pi/2\right)$ is used to switch $Y$ and $Z$ errors. Since this protocol can detect $X$, $Y$, and $Z$ errors over two sequential iterations, it can detect any single-qubit error according to the discretization of quantum errors [91]. This explains why the recurrence protocols are effective in various error models, not just the Werner-state model.
Based on the above insight, we can extend the $2$-to-$1$ recurrence protocol to a higher-rate $n$-to-$\left(n-1\right)$ protocol using a stabilizer code $X_{1}X_{2}\cdots X_{n}$. For example, consider the $3$-to-$2$ entanglement distillation circuit illustrated in Figure  2, constructed using the unitary decoding method. The two dashed boxes represent circuits for two sequential iterations. In each box, the projective measurement on the first qubit reveals the parity of $X_{1}X_{2}X_{3}$. By comparing parities, Alice and Bob can detect single-qubit $Y$ or $Z$ errors. Hadamard gates on the remaining qubits convert $X$ errors to $Z$ errors, which are detected in the next iteration. The protocol is similar to the generalized Shor’s construction of a quantum code [92], with one code addressing bit-flip errors and the other addressing phase-flip errors. By permuting the measurements and gates in Figure 2, we obtain the stabilizers for two iterations:
@@
which characterize a $\left[\!\left[9,4,2\right]\!\right]$ quantum code capable of detecting a single-qubit error. For the $n$-to-$\left(n-1\right)$ protocol, two iterations correspond to a $\left[\!\left[n^{2},\left(n-1\right)^{2},2\right]\!\right]$ quantum code. The term “$n$-to-$(n-1)$ protocol” reflects that each iteration can be executed independently, requiring only $n$ inputs, corresponding to a classical $\left[n,n-1,2\right]$ error-detecting code. It corrects bit-flip or phase-flip errors, effectively improving output fidelity similarly to the original $2$-to-$1$ protocol.
To prevent correlated multiqubit errors, it is essential to use outputs from different iterations as inputs in subsequent iterations. For example, in Figure 2, an $X_{1}$ error on the control qubit can propagate through the CNOT gates to become $X_{1}X_{2}X_{3}$. Assuming a depolarizing channel with an error probability of $p$ per input qubit, the effective error probability in the first iteration is $\frac{2p}{3}$ due to $Y$ and $Z$ errors. In the second iteration, the effective error probability is also $\frac{2p}{3}$ due to undetected $X$ errors on the output and control qubits. Thus, if the input meets the fidelity threshold initially, it will meet the threshold in subsequent iterations. This result applies to other $n$-to-$\left(n-1\right)$ recurrence protocols as well.
@@
where $k$ and $n$ are the logical and physical qubits in an $\left[\!\left[n,k,d\right]\!\right]$ code, and $P_{S}$ is the success probability. Our simulation uses the QuantumClifford.jl package, which utilizes the stabilizer tableaux formalism [93]. To evaluate the performance of protocols capable of correcting all types of quantum errors, we simulate two sequential iterations per protocol, corresponding to an $n^{2}$-to-$\left(n-1\right)^{2}$ entanglement distillation that corrects both bit-flip and phase flip errors. We use Werner states as inputs, which are mixtures of a standard Bell state and a completely mixed state, as described by Equation 2. We assume error-free local quantum operations and classical communication, a valid assumption given that channel noise significantly exceeds local noise.
@@
Figure 3(a) shows the fidelity of distilled Bell states as a function of input fidelity for several distillation protocols. Circles mark the threshold where output fidelity exceeds input fidelity, which increases with the code rate. Figure 3(b) shows yield as a function of input fidelity. The $4$-to-$1$ protocol has the highest yield for input fidelities from $0.5$ to $0.887$. Beyond this range, higher-rate protocols are superior. Near ideal fidelity, yield is determined by code rate. Achieving optimal yield at high fidelity requires large $n$ in $n^{2}$-to-$\left(n-1\right)^{2}$ protocols. In Appendix B, we demonstrate that adaptively increasing the block size $n$ based on input fidelity allows for a constant overall yield rate for any desired fidelity.
@@
We can further extend the recurrence protocol by adapting classical codes. In the first iteration, the code corrects bit-flip errors. Hadamard gates then convert the remaining phase-flip errors to bit-flip errors, which the code can correct in the second iteration. Since these are classical codes, the commutation condition for each stabilizer is automatically met. Moreover, each iteration is flexible in its choice of classical code, allowing customization based on the strength of phase-flip and bit-flip errors. For example, if phase-flip errors dominate, we can apply a high-rate classical code in the first (bit-flip) iteration and a lower-rate code with greater error-correcting capability in the second (phase-flip) iteration. This recurrence entanglement distillation supports the direct implementation of well-developed classical error-correcting codes, such as polar codes, LDPC codes, and turbo codes. Although the generalized Shor construction can build quantum codes [92], this method sacrifices certain desired properties. For example, combining two classical LDPC codes via Shor’s construction does not produce a quantum LDPC code [94]. In contrast, our approach can apply classical LDPC codes in separate iterations, preserving the sparsity of parity-check matrices. While quantum LDPC codes can be constructed using the hypergraph product [94], the recurrence protocol simplifies the direct use of classical LDPC codes and can readily extend to other types of codes. In practice, noisy Bell states can be non-Werner [95, 96, 97]. Combining the stabilizer formalism with fidelity and yield aids in designing effective distillation protocols.
IV Decoding the Stabilizer Code by Single-Qubit Measurements